At 400 K,the root mean square(rms) speed of a gas X (molecular weight = 40) is equal to the most probable speed of gas Y at 60 K. The molecular weight of the gas Y is.
 

Grade:11

3 Answers

Khimraj
5 years ago
Urms = Ump
\sqrt{3RT/M_{X}} =\sqrt{2RT/M_{Y}}
\sqrt{3R*400/40} =\sqrt{2R*60/M_{Y}}
MY = 4
Hope it clears.
Arun
5 years ago
Dear student
 
 
hope it helps
In case of any query, please feel free to ask.
 
Regards
Arun
Arun
5 years ago
Dear student
 
Vrms = Vmps
 
sqrt (3R*400/40) = sqrt (2R *60 / M)
 
30 R = 120 R/M
 
M = 4
 
Regards
Arun